Massachusetts Institute of Technology Department of Electrical Engineering and Computer Science

Proposal for Thesis Research in Partial Fulfillment of the Requirements for the Degree of Master of Science

Title: Submitted by: (name

(address)  (Signature of author)

Date of submission: Expected Date of Completion: Laboratory where thesis will be done: Brief Statement of the Problem: Supervision Agreement: The program outlined in this proposal is adequate for a Master’s thesis. The supplies and facilities required are available, and I am willing to supervise the research and evaluate the thesis report.